How much money does Jessica have in her Education Savings plan if, starting today, she can withdraw $350 per month for four years and her plan earns 7.2% compounded monthly?

Select one:

a. $16,800

b. $10,782

c. $14,647

d. $11,962

e. $13,494
Answer #1

Solution

Answer-$14647

This is an annuity due as the first payment is made at the beginning of period

Present value of annuity due=Annuity payment*((1-(1/(1+r)^n))/r)*(1+r)

Where n=number of periods=12*4=48

r=rate of intrest per period=7.2/12=.6% per month

Annuity payment=350

Present value of annuity due=350*((1-(1/(1+.006)^48))/.006)*(1+.006)

=14646.94426

Thus correct answer is 14647

If you are satisfied with the answer,please give a thumbs up